OpenStudy (3psilon):

SAT help ? In the equation, K and m are constants . If the equation is true for all values of x, what is the value of m?

OpenStudy (3psilon):

\[(x-8)(x-k) = x^{2} -5kx + m\]

OpenStudy (bahrom7893):

simplify the left side: x^2 - kx - 8x + 8k = x^2 - (k+8)x + 8k

OpenStudy (bahrom7893):

x^2 - (k+8)x + 8k = x^2 - 5kx + m

OpenStudy (bahrom7893):

thus you know that k+8 = 5k

OpenStudy (bahrom7893):

and 8k = m

OpenStudy (anonymous):

just what bahrom said

OpenStudy (bahrom7893):

Now just solve for both.

OpenStudy (anonymous):

bahrom mate can ya help me out with a sum of my own??

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

@robinfr93 let me see the question

OpenStudy (3psilon):

@bahrom7893 What does it mean by true for all values of x?

OpenStudy (3psilon):

I don't see how you got k+8 = 5k Where did everything else go?

OpenStudy (bahrom7893):

it just means equate the two equations.

OpenStudy (bahrom7893):

I'm equating the coefficients: x^2 - (k+8)x + 8k = x^2 - 5kx + m

OpenStudy (bahrom7893):